OpenStudy (thefluffmuncher):

Simplify. squareroot of 64 A.–4 B.4 C.–8 D.8

OpenStudy (anonymous):

I'll teach you this

OpenStudy (anonymous):

First of all it is 8

OpenStudy (anonymous):

8 but you could just use a calculator

OpenStudy (thefluffmuncher):

ok

OpenStudy (anonymous):

8*8=64

OpenStudy (thefluffmuncher):

yea

OpenStudy (thefluffmuncher):

so its 8

OpenStudy (anonymous):

Yes, it's 8.. The square root of 25 is 5, the square root of 16 is 4, the square root of 100, is 10

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

See the pattern, 5*5=25,6*6=36,10*10=100

OpenStudy (thefluffmuncher):

yay ty so much